OpenStudy (anonymous):

3root4-2root14+12-root14 I've have tried to simplify this, but I still don't understand this. Please help!

OpenStudy (anonymous):

The answer is supposed to be 18-3root14

OpenStudy (ghazi):

\[3\sqrt{4}-2\sqrt{14}+12\sqrt{14}= 3*2+10\sqrt{14}=6+10\sqrt{14}=2(3+5\sqrt{14})\]

OpenStudy (anonymous):

no, you messed up because its 12-root14 not 12root14, so the answer turns out wrong

OpenStudy (ghazi):

oh...i see

OpenStudy (ghazi):

still you can do it...it's 18-3 root 14

OpenStudy (ghazi):

\[3\sqrt{4}-2\sqrt{14}+12-\sqrt{14}=12+6-3\sqrt{14}=18-3\sqrt{14}\]

OpenStudy (ghazi):

@nppbleh789 is it clear?

OpenStudy (anonymous):

ohh, i get it. Thankyou (:

OpenStudy (ghazi):

:)